# An Attack on $p$-adic Lattice Public-key Cryptosystems and Signature Schemes

### v1

@@ Abstract
2010 Mathematics Subject Classification: Primary 11F85, Secondary 94A60.
Key words and phrases: $p$-adic lattice, LVP, CVP, Signature scheme, Public-key cryptosystem, Attack.
@@ 1 Introduction
Lattices were investigated since the late $18$th century. They are important research objects in geometric number theory and have many applications. There are two famous computational problems in Euclidean lattices: the Shortest Vector Problem (SVP) and the Closest Vector Problem (CVP). Many researches have been done to study these problems. Van Emde Boas [10] proved that the SVP is NP-hard in the $l_{\infty}$ norm and the CVP is NP-hard in the $l_{2}$ norm by reducing them to the weak partition problem. Ajtai [11] proved that the SVP is NP-hard under randomized reductions in the $l_{2}$ norm. These hard problems are useful in constructing cryptographic primitives. The GGH scheme [12], the NTRU scheme [13] and the LWE scheme [14] are based on these problems.
@@ 1.1 Previous Works
The experimental results demonstrated that the new schemes deserve good efficiency. As for security, there is no formal proof for these schemes since the problems in local fields are relatively very new. Instead, Deng et al. [3] discussed some possible attacks on their schemes, including recovering a uniformizer, finding an orthogonal basis, solving CVP-instances and modulo $p$ attack. Therefore, the security of their schemes remains an open problem.
1.2 Our Contribution
In order to figure out whether these schemes are secure, we have made tremendous efforts. Unfortunately, these schemes turn out to be insecure. There are efficient algorithms to forge a valid signature of any message and decrypt any ciphertext with the public key. The basic principle of the attack is that it is easy to solve the LVP and the CVP in totally ramified extension fields when $p$ is a polynomial in the rank of the given lattice.
Consider a totally ramified extension field $K$ over $\mathbb{Q}_{p}$ of degree $n$, and let $\pi$ be a uniformizer of $K$. Then $\left|\pi\right|_{p}=p^{-\frac{1}{n}}$ and the ring $\mathcal{O}_{K}$ of algebraic integers of $K$ is $\mathbb{Z}_{p}[\pi]$. Each vector $\alpha\in\mathcal{O}_{K}$ can be written as
$\alpha=a_{0}+a_{1}\pi+a_{2}\pi^{2}+\cdots+a_{n-1}\pi^{n-1},$
where $a_{0},a_{1},\dots,a_{n-1}\in\mathbb{Z}_{p}$. There is only one summand that contributes to the $p$-adic absolute value of $\alpha$, i.e., $\left|\alpha\right|_{p}=\left|a_{i}\pi^{i}\right|_{p}$ for some $10\leq i\leq n-$ 1. Therefore, if two vectors $\alpha,\beta\in\mathcal{O}_{K}$ have the same $p$-adic absolute value, then there must be a strictly shorter vector $\alpha-c\beta$ for a certain $c\in\{0,1,\dots,p-1\}$. For example, suppose that the constant coefficient of $\alpha$ and $\beta$ is $a_{0}$ and $b_{0}$ respectively, both in $\mathbb{Z}_{p}-p\mathbb{Z}_{p}$, then $c=a_{0}b_{0}^{-1}\mod p$. In general, since the uniformizer $\pi$ is secret, we need to compute all $p$ possibilities to obtain this short vector.
Based on this result, when $p$ is a polynomial in $n$, we can solve the LVP in the lattice $\mathcal{L}(1,\pi,\pi^{2},\dots,\pi^{n-1})=\mathbb{Z}_{p}[\pi]$ efficiently even if the lattice is written in another basis which hides the uniformizer. The LVP algorithm takes $O(pn)$ many $p$-adic absolute value computations of elements of $K$. Subsequently, we can recover a uniformizer and an orthogonal basis of this lattice, which enables us to forge a valid signature of any message and decrypt any ciphertext. Hence the public-key encryption cryptosystem and signature scheme are broken completely. The attack takes $O(pn^{2})$ many $p$-adic absolute value computations of elements of $K$.
Nevertheless, it does not mean that $p$-adic lattices are not suitable in constructing cryptographic primitives. In a general extension field $K$ over $\mathbb{Q}_{p}$, the degree $n$ equals $ef$ where $e$ is the ramification index and $f$ is the residue degree. The above LVP algorithm takes $O(mp^{f})$ many $p$-adic absolute value computations of elements of $K$, where $m$ is the rank of the input lattice. Comparing with the LVP algorithm given in [1, 2], which takes $O(p^{m})$ many $p$-adic absolute value computations, it is a considerable improvement. However, it is still an exponential time algorithm even though $p$ is a polynomial in $m$. Our attack is inefficient when the extension field has a large residue degree. Therefore, the schemes can be modified to resist our attack.
Paper Organization
@@ 2.1 Basic Facts about Local Fields
the residue field degree of $K/\mathbb{Q}_{p}$. As ideals in $\mathcal{O}_{K}$, we have $p\mathcal{O}_{K}=\pi^{e}\mathcal{O}_{K}$. Call the positive integer $e$ the ramification index of $K/\mathbb{Q}_{p}$. We have $n=\left[K:\mathbb{Q}_{p}\right]=ef$. When $e=1$, the extension $K/\mathbb{Q}_{p}$ is unramified, and when $e=n$, $K/\mathbb{Q}_{p}$ is totally ramified. Each element $x$ of the multiplicative group $K^{\times}$ of nonzero elements of $K$ can be written uniquely as $x=u\pi^{t}$ with $u\in\mathcal{O}_{K}^{\times}$ and $t\in\mathbb{Z}$. We have $p=u\pi^{e}$ with $u\in\mathcal{O}_{K}^{\times}$, so $\left|\pi\right|_{p}=p^{-\frac{1}{e}}$. The valuation group of $K$ is
$\left\{\left|x\right|_{p}:x\in K^{\times}\right\}=p^{\frac{\mathbb{Z}}{e}}.$
@@ 2.2 Norm and Orthogonal Basis
Let $p$ be a prime. Let $V$ be a left vector space over $\mathbb{Q}_{p}$. A norm on $V$ is a function
@@
Definition 2.2 (orthogonal basis [8]).
Let $V$ be a left vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. We call $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ an orthogonal basis of $V$ over $\mathbb{Q}_{p}$ if $V$ can be decomposed into the direct sum of $n$ $1$-dimensional subspaces $V_{i}$’s $(1\leq i\leq n)$, such that
@@
for any $v_{i}\in V_{i}$, $i=1,2,\dots,n$, where $V_{i}$ is spanned by $\alpha_{i}$. Two subspaces $U$, $W$ of $V$ is said to be orthogonal if the sum $U+W$ is a direct sum and it holds that $\left|u+w\right|=\max\left\{\left|u\right|,\left|w\right|\right\}$ for all $u\in U$, $w\in W$.
@@
Let $V$ be a left vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be a basis of $V$ over $\mathbb{Q}_{p}$. If
$\left\{\left|v_{i}\right|:v_{i}\in\mathbb{Q}_{p}\cdot\alpha_{i}\right\}\bigcap\left\{\left|v_{j}\right|:v_{j}\in\mathbb{Q}_{p}\cdot\alpha_{i}\right\}=\left\{0\right\}$
@@
Let $V$ be a left vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be $\mathbb{Q}_{p}$-linearly independent vectors of $V$. Then $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ is an orthogonal basis of $V$ if and only if it holds that
@@ 2.3 $p$-adic Lattice
Definition 2.5 ($p$-adic lattice [1, 2]).
Let $p$ be a prime number, and let $K$ be an extension field of $\mathbb{Q}_{p}$ of degree $n$, where $n$ is a positive integer. Let $m$ be a positive integer with $1\leq m\leq n$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{m}\in K$ be $\mathbb{Q}_{p}$-linearly independent vectors. A $p$-adic lattice in $K$ is the set
$\mathcal{L}(\alpha_{1},\alpha_{2},\dots,\alpha_{m}):=\left\{\sum^{m}_{i=1}{a_{i}\alpha_{i}}:a_{i}\in\mathbb{Z}_{p},1\leq i\leq m\right\}$
of all $\mathbb{Z}_{p}$-linear combinations of $\alpha_{1},\alpha_{2},\dots,\alpha_{m}$. The sequence of vectors $\alpha_{1},\alpha_{2},\dots,\alpha_{m}$ is called a basis of the lattice $\mathcal{L}(\alpha_{1},\alpha_{2},\dots,\alpha_{m})$. The integers $m$ and $n$ are called the rank and dimension of the lattice, respectively. When $n=m$, we say that the lattice is of full rank.
@@
Definition 2.6 (orthogonal basis of a $p$-adic lattice [4]).
[…]
@@ 2.5 Successive Maxima
Definition 2.10 (successive maxima [5]).
Let $V$ be a left vector space over $\mathbb{Q}_{p}$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\mathcal{L}$ be a lattice of rank $n$ in $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be an orthogonal basis of $\mathcal{L}$ such that $\left|\alpha_{1}\right|\geq\left|\alpha_{2}\right|\geq\cdots\geq\left|\alpha_{n}\right|$. The $i$th successive maxima of $\mathcal{L}$ respect to norm ||$\cdot$|| ⋅ | is
@@
Let $V$ be a left vector space over $\mathbb{Q}_{p}$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\mathcal{L}$ be a lattice of rank $n$ in $V$. Suppose that $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ and $\beta_{1},\beta_{2},\dots,\beta_{n}$ are two orthogonal bases of $\mathcal{L}$ such that $\left|\alpha_{1}\right|\geq\left|\alpha_{2}\right|\geq\cdots\geq\left|\alpha_{n}\right|$ and $\left|\beta_{1}\right|\geq\left|\beta_{2}\right|\geq\cdots\geq\left|\beta_{n}\right|$. Then we have $\left|\alpha_{i}\right|=\left|\beta_{i}\right|$ for $1\leq i\leq n$.
@@ 2.6.1 The $p$-adic Signature Scheme
The correctness is obvious. For the efficiency, we can prove that for a fixed $f(x)$, the probability of $\mathbb{Z}_{p}[\theta]=\mathbb{Z}_{p}[\zeta]$ is $1-\frac{1}{p}$. Also, we can always get a valid signature with just one $r$. See the following lemmas.
Lemma 2.12.
Write
$\zeta=a_{0}+a_{1}\theta+a_{2}\theta^{2}+\cdots+a_{n-1}\theta^{n-1}$
Proof.
$\mathbb{Z}_{p}[\zeta]=\mathbb{Z}_{p}[\zeta-a_{0}]=\mathbb{Z}_{p}[\theta].$
Lemma 2.13.
For any $t=H(M\|r)$, we have $\left|t-v\right|_{p}<1$.
Proof.
Since $K=\mathbb{Q}_{p}(\theta)$, we can write
$t=a_{1}\alpha_{1}+a_{2}\alpha_{2}+\cdots+a_{n}\alpha_{n}$
@@ 3 Orthogonalization in Local Fields
As mentioned in [3] and [5], the LVP and the CVP can be solved in polynomial time with the help of an orthogonal bases. Therefore, we consider orthogonalization algorithms straightly. In totally ramified fields, the simplified orthogonalization algorithm takes $O(pm^{2})$ many $p$-adic absolute value computations of elements of $K$, where $m$ is the rank of the input lattice. It is a polynomial time algorithm when $p$ is a polynomial in $m$.
@@ 3.1 Orthogonalization in Totally Ramified Fields
By Proposition 2.3, $1,\pi,\pi^{2},\dots,\pi^{n-1}$ is an orthogonal basis of $\mathcal{O}_{K}$. The crucial point is that this orthogonal basis satisfies
@@
By Proposition 2.3, $\alpha_{1},\alpha_{2},\dots,\alpha_{m}$ is an orthogonal basis of $\mathcal{L}$. Let $A=(a_{ij})\in{\rm GL}_{m}(\mathbb{Z}_{p})$ be an invertible matrix over $\mathbb{Z}_{p}$ such that
@@
$\begin{array}[]{c}\beta_{1}=a_{11}\alpha_{1}+a_{12}\alpha_{2}+\cdots+a_{1m}\alpha_{m}\\ \beta_{2}=a_{21}\alpha_{1}+a_{22}\alpha_{2}+\cdots+a_{2m}\alpha_{m}\\ \vdots\\ \beta_{m}=a_{m1}\alpha_{1}+a_{12}\alpha_{2}+\cdots+a_{mm}\alpha_{m}.\end{array}$
If $a_{i1}\in\mathbb{Z}_{p}-p\mathbb{Z}_{p}$, then we have $\left|\beta_{i}\right|_{p}=\left|\alpha_{1}\right|_{p}=\lambda_{1}$. If $a_{i1}\in p\mathbb{Z}_{p}$, then we have $\left|\beta_{i}\right|_{p}\leq\left|\alpha_{2}\right|_{p}<\lambda_{1}$. We may assume that for $1\leq i\leq l$ $(l\geq 2)$, $a_{i1}\in\mathbb{Z}_{p}-p\mathbb{Z}_{p}$ and for $l+1\leq i\leq m$, $a_{i1}\in p\mathbb{Z}_{p}$. Then for each $2\leq i\leq l$, there is an integer $c_{i}\in\{0,1,\dots,p-1\}$ such that $a_{i1}-c_{i}a_{11}\in p\mathbb{Z}_{p}$ and hence
@@
$\begin{array}[]{c}\beta_{1}^{\prime}=a_{11}^{\prime}\alpha_{1}+a_{12}^{\prime}\alpha_{2}+\cdots+a_{1m}^{\prime}\alpha_{m}\\ \beta_{2}^{\prime}=a_{21}^{\prime}\alpha_{1}+a_{22}^{\prime}\alpha_{2}+\cdots+a_{2m}^{\prime}\alpha_{m}\\ \vdots\\ \beta_{m}^{\prime}=a_{m1}^{\prime}\alpha_{1}+a_{12}^{\prime}\alpha_{2}+\cdots+a_{mm}^{\prime}\alpha_{m},\end{array}$
@@
Algorithm (finding $\lambda_{2}$).
Input: a lattice $\mathcal{L}=\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$.
Output: the value $\lambda_{2}$ and a lattice vector $v\in\mathcal{L}$ such that $\left|v\right|_{p}=\lambda_{2}$.
  1. 1.
    rearrange $\beta_{1},\beta_{2},\dots,\beta_{m}$ such that $\left|\beta_{1}\right|_{p}=\max{\left\{\left|\beta_{1}\right|_{p},\left|\beta_{2}\right|_{p},\dots,\left|\beta_{m}\right|_{p}\right\}}$,
  2. 2.
    $\beta_{1}^{\prime}\leftarrow\beta_{1}$,
  3. 3.
    for $i=2$ to $m$ do:
  4. 4.
    for $j=0$ to $p-1$ do:
  5. 5.
    if $\left|\beta_{i}-j\beta_{1}\right|_{p}<\left|\beta_{1}\right|_{p}$ then $\beta_{i}^{\prime}\leftarrow\beta_{i}-j\beta_{1}$,
  6. 6.
    let $s$ be the subscript such that $\left|\beta_{s}^{\prime}\right|_{p}=\max{\left\{\left|\beta_{2}^{\prime}\right|_{p},\left|\beta_{3}^{\prime}\right|_{p},\dots,\left|\beta_{m}^{\prime}\right|_{p}\right\}},$
Return $\lambda_{2}\leftarrow\left|\beta_{s}^{\prime}\right|_{p}$, $v\leftarrow\beta_{s}^{\prime}$.
@@
Let $A$ be an algorithm having the same input and procedure as the algorithm “finding $\lambda_{2}$”, but outputting the new basis, i.e., $A(\beta_{1},\beta_{2},\dots,\beta_{m})=\beta_{1}^{\prime},\beta_{2}^{\prime},\dots,\beta_{m}^{\prime}$. We summarize the algorithm in above theorem as follows.
Algorithm (finding an orthogonal basis).
Input: a lattice $\mathcal{L}=\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$.
Output: an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$.
  1. 1.
    $\beta_{1}^{\prime},\beta_{2}^{\prime},\dots,\beta_{m}^{\prime}\leftarrow\beta_{1},\beta_{2},\dots,\beta_{m}$,
  2. 2.
    for $i=1$ to $m-1$ do:
  3. 3.
    $\beta_{i}^{\prime},\beta_{i+1}^{\prime},\dots,\beta_{m}^{\prime}\leftarrow A(\beta_{i}^{\prime},\beta_{i+1}^{\prime},\dots,\beta_{m}^{\prime})$,
Return $\beta_{1}^{\prime},\beta_{2}^{\prime},\dots,\beta_{m}^{\prime}$.

### v2

# An Attack on $p$-adic Lattice Public-key Encryption Cryptosystems and Signature Schemes

@@ Abstract
keywords:
$p$-adic lattice, LVP, CVP, Signature scheme, Public-key encryption cryptosystem, Attack
@@ 1 Introduction
Lattices were investigated since the late $18$th century. They are important research objects in geometric number theory and have many applications. There are two famous computational problems in Euclidean lattices: the Shortest Vector Problem (SVP) and the Closest Vector Problem (CVP). Much research has been done to study these problems. Van Emde Boas [11] proved that the SVP is NP-hard in the $l_{\infty}$ norm and the CVP is NP-hard in the $l_{2}$ norm by reducing them to the weak partition problem. Ajtai [12] proved that the SVP is NP-hard under randomized reductions in the $l_{2}$ norm. These hard problems are useful in constructing cryptographic primitives. The GGH scheme [13], the NTRU scheme [14] and the LWE scheme [15] are based on these problems.
@@
The experimental results demonstrated that the new schemes achieve good efficiency. As for security, there is no formal proof for these schemes since the problems in local fields are relatively very new. Instead, Deng et al. [3] discussed some possible attacks on their schemes, including recovering a uniformizer, finding an orthogonal basis, solving CVP-instances and modulo $p$ attack. Therefore, the security of their schemes remains an open problem.
@@ 2.1 Basic Facts about Local Fields
the residue field degree of $K/\mathbb{Q}_{p}$. As ideals in $\mathcal{O}_{K}$, we have $p\mathcal{O}_{K}=\pi^{e}\mathcal{O}_{K}$. Call the positive integer $e$ the ramification index of $K/\mathbb{Q}_{p}$. We have $n=\left[K:\mathbb{Q}_{p}\right]=ef$. When $e=1$, the extension $K/\mathbb{Q}_{p}$ is unramified, and when $e=n$, $K/\mathbb{Q}_{p}$ is totally ramified. Each element $x$ of the multiplicative group $K^{\times}$ of nonzero elements of $K$ can be written uniquely as $x=u\pi^{t}$ with $u\in\mathcal{O}_{K}^{\times}$ and $t\in\mathbb{Z}$. We have $p=u\pi^{e}$ with $u\in\mathcal{O}_{K}^{\times}$, so $\left|\pi\right|_{p}=p^{-\frac{1}{e}}$.
@@ 2.2 Norm and Orthogonal Basis
Let $p$ be a prime. Let $V$ be a vector space over $\mathbb{Q}_{p}$. A norm on $V$ is a function
@@
Definition 1 (orthogonal basis [9]).
Let $V$ be a vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. We call $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ an orthogonal basis of $V$ over $\mathbb{Q}_{p}$ if $V$ can be decomposed into the direct sum of $n$ $1$-dimensional subspaces $V_{i}={\rm Span}_{\mathbb{Q}_{p}}(\alpha_{i})$ $(1\leq i\leq n)$, such that
@@
for any $v_{i}\in V_{i}$, $i=1,2,\dots,n$. Two subspaces $U$, $W$ of $V$ is said to be orthogonal if the sum $U+W$ is a direct sum and it holds that $\left|u+w\right|=\max\left\{\left|u\right|,\left|w\right|\right\}$ for all $u\in U$, $w\in W$.
@@
Let $V$ be a vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be a basis of $V$ over $\mathbb{Q}_{p}$. If
$\left\{\left|v_{i}\right|:v_{i}\in\mathbb{Q}_{p}\cdot\alpha_{i}\right\}\bigcap\left\{\left|v_{j}\right|:v_{j}\in\mathbb{Q}_{p}\cdot\alpha_{j}\right\}=\left\{0\right\}$
@@
Let $V$ be a vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be $\mathbb{Q}_{p}$-linearly independent vectors of $V$. Then $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ is an orthogonal basis of $V$ if and only if it holds that
@@ 2.3 $p$-adic Lattice
Definition 2 ($p$-adic lattice [4]).
Let $V$ be a vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$. Let $m$ be a positive integer with $1\leq m\leq n$. Let $\alpha_{1},\dots,\alpha_{m}\in V$ be $\mathbb{Q}_{p}$-linearly independent vectors. A $p$-adic lattice in $V$ is the set
$\mathcal{L}(\alpha_{1},\dots,\alpha_{m}):=\left\{\sum^{m}_{i=1}{a_{i}\alpha_{i}}:a_{i}\in\mathbb{Z}_{p},1\leq i\leq m\right\}$
of all $\mathbb{Z}_{p}$-linear combinations of $\alpha_{1},\dots,\alpha_{m}$. The sequence of vectors $\alpha_{1},\dots,\alpha_{m}$ is called a basis of the lattice $\mathcal{L}(\alpha_{1},\dots,\alpha_{m})$. The integers $m$ and $n$ are called the rank and dimension of the lattice, respectively. When $n=m$, we say that the lattice is of full rank.
@@
Definition 3 (orthogonal basis of a $p$-adic lattice [4]).
Let $V$ be nontrivial finite-dimensional vector space over $\mathbb{Q}_{p}$, and let $\alpha_{1},\dots,\alpha_{n}\in V$. If $\alpha_{1},\dots,\alpha_{n}\in V$ form an orthogonal basis of the vector space ${\rm Span_{\mathbb{Q}_{p}}}(\alpha_{1},\dots,\alpha_{n})$, then we say $\alpha_{1},\dots,\alpha_{n}$ is an orthogonal basis of the lattice $\mathcal{L}(\alpha_{1},\dots,\alpha_{n})$.
@@ 2.5 Successive Maxima
Definition 7 (successive maxima [5]).
Let $V$ be a vector space over $\mathbb{Q}_{p}$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\mathcal{L}$ be a lattice of rank $n$ in $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be an orthogonal basis of $\mathcal{L}$ such that $\left|\alpha_{1}\right|\geq\left|\alpha_{2}\right|\geq\cdots\geq\left|\alpha_{n}\right|$. The $i$th successive maxima of $\mathcal{L}$ respect to norm ||$\cdot$|| ⋅ | is
@@
Let $V$ be a vector space over $\mathbb{Q}_{p}$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\mathcal{L}$ be a lattice of rank $n$ in $V$. Suppose that $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ and $\beta_{1},\beta_{2},\dots,\beta_{n}$ are two orthogonal bases of $\mathcal{L}$ such that $\left|\alpha_{1}\right|\geq\left|\alpha_{2}\right|\geq\cdots\geq\left|\alpha_{n}\right|$ and $\left|\beta_{1}\right|\geq\left|\beta_{2}\right|\geq\cdots\geq\left|\beta_{n}\right|$. Then we have $\left|\alpha_{i}\right|=\left|\beta_{i}\right|$ for $1\leq i\leq n$.
@@ 2.6.1 The $p$-adic Signature Scheme
The correctness is obvious. For the efficiency, we can prove that for a fixed $f(x)$, the probability of $\mathbb{Z}_{p}[\theta]=\mathbb{Z}_{p}[\zeta]$ is $1-\frac{1}{p}$. Also, we can always get a valid signature with just one $r$. See Lemma A.1 and Lemma A.2 in Appendix A.
@@ 3 Orthogonalization in Local Fields
As mentioned in [3] and [5], the LVP and the CVP can be solved in polynomial time with the help of an orthogonal basis. Therefore, we consider orthogonalization algorithms directly. In totally ramified fields, the simplified orthogonalization algorithm takes $O(pm^{2})$ many $p$-adic absolute value computations of elements of $K$, where $m$ is the rank of the input lattice. It is a polynomial time algorithm when $p$ is a polynomial in $m$.
@@ 3.1 Orthogonalization in Totally Ramified Fields
By Proposition 3.4 in [3], $1,\pi,\pi^{2},\dots,\pi^{n-1}$ is an orthogonal basis of $\mathcal{O}_{K}$. The crucial point is that this orthogonal basis satisfies
@@
Since
$\left\{\left|v_{i}\right|:v_{i}\in\mathbb{Q}_{p}\cdot\alpha_{i}\right\}=\{0\}\cup p^{\mathbb{Z}}\cdot\left|\alpha_{i}\right|$
and
$\left|\alpha_{1}\right|_{p}>\left|\alpha_{2}\right|_{p}>\cdots>\left|\alpha_{m}\right|_{p}>\left|p\alpha_{1}\right|_{p},$
we have
$\left\{\left|v_{i}\right|:v_{i}\in\mathbb{Q}_{p}\cdot\alpha_{i}\right\}\bigcap\left\{\left|v_{j}\right|:v_{j}\in\mathbb{Q}_{p}\cdot\alpha_{j}\right\}=\left\{0\right\}$
for any $i,j=1,2,\dots,n$, $i\neq j$. Therefore, by Proposition 2.2, $\alpha_{1},\alpha_{2},\dots,\alpha_{m}$ is an orthogonal basis of $\mathcal{L}$. Let $A=(a_{ij})\in{\rm GL}_{m}(\mathbb{Z}_{p})$ be an invertible matrix over $\mathbb{Z}_{p}$ such that
@@
$\begin{array}[]{c}\beta_{1}=a_{11}\alpha_{1}+a_{12}\alpha_{2}+\cdots+a_{1m}\alpha_{m}\\ \beta_{2}=a_{21}\alpha_{1}+a_{22}\alpha_{2}+\cdots+a_{2m}\alpha_{m}\\ \vdots\\ \beta_{m}=a_{m1}\alpha_{1}+a_{m2}\alpha_{2}+\cdots+a_{mm}\alpha_{m}.\end{array}$
If $a_{i1}\in\mathbb{Z}_{p}-p\mathbb{Z}_{p}$, then we have $\left|\beta_{i}\right|_{p}=\left|\alpha_{1}\right|_{p}=\lambda_{1}$. If $a_{i1}\in p\mathbb{Z}_{p}$, then we have $\left|\beta_{i}\right|_{p}\leq\left|\alpha_{2}\right|_{p}<\lambda_{1}$. Since the longest vector of $\mathcal{L}$ is $\alpha_{1}$, there must be at least one $\beta_{i}$ such that $\left|\beta_{i}\right|_{p}=\lambda_{1}$. Hence, we may assume that for $1\leq i\leq l$, $a_{i1}\in\mathbb{Z}_{p}-p\mathbb{Z}_{p}$ and for $l+1\leq i\leq m$, $a_{i1}\in p\mathbb{Z}_{p}$. If $l=1$, then
$\lambda_{2}=\max{\left\{\left|\beta_{2}\right|_{p},\left|\beta_{3}\right|_{p},\dots,\left|\beta_{m}\right|_{p}\right\}}.$
We can find the second longest vector $v$ directly from $\beta_{2},\beta_{3},\dots,\beta_{m}$. Now we assume that $l\geq 2$. Then for each $2\leq i\leq l$, there is an integer $c_{i}\in\{0,1,\dots,p-1\}$ such that $a_{i1}-c_{i}a_{11}\in p\mathbb{Z}_{p}$ and hence
@@
$\begin{array}[]{c}\beta_{1}^{\prime}=a_{11}^{\prime}\alpha_{1}+a_{12}^{\prime}\alpha_{2}+\cdots+a_{1m}^{\prime}\alpha_{m}\\ \beta_{2}^{\prime}=a_{21}^{\prime}\alpha_{1}+a_{22}^{\prime}\alpha_{2}+\cdots+a_{2m}^{\prime}\alpha_{m}\\ \vdots\\ \beta_{m}^{\prime}=a_{m1}^{\prime}\alpha_{1}+a_{m2}^{\prime}\alpha_{2}+\cdots+a_{mm}^{\prime}\alpha_{m},\end{array}$
@@
Algorithm 1 finding $\lambda_{2}$
1:a lattice $\mathcal{L}=\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$
2:the value $\lambda_{2}$ and a lattice vector $v\in\mathcal{L}$ such that $\left|v\right|_{p}=\lambda_{2}$
3:rearrange $\beta_{1},\beta_{2},\dots,\beta_{m}$ such that $\left|\beta_{1}\right|_{p}=\max{\left\{\left|\beta_{1}\right|_{p},\left|\beta_{2}\right|_{p},\dots,\left|\beta_{m}\right|_{p}\right\}}$
4:$\beta_{1}^{\prime}\Leftarrow\beta_{1}$
5:for $i=2$ to $m$ do
6:     for $j=0$ to $p-1$ do
7:         if $\left|\beta_{i}-j\beta_{1}\right|_{p}<\left|\beta_{1}\right|_{p}$ then
8:              $\beta_{i}^{\prime}\Leftarrow\beta_{i}-j\beta_{1}$
9:         end if
10:     end for
11:end for
12:let $s$ be the subscript such that $\left|\beta_{s}^{\prime}\right|_{p}=\max{\left\{\left|\beta_{2}^{\prime}\right|_{p},\left|\beta_{3}^{\prime}\right|_{p},\dots,\left|\beta_{m}^{\prime}\right|_{p}\right\}}$
13:return $\lambda_{2}\Leftarrow\left|\beta_{s}^{\prime}\right|_{p}$, $v\Leftarrow\beta_{s}^{\prime}$
@@
Let $A$ be an algorithm having the same input and procedure as Algorithm 1, but outputting the new basis, i.e., $A(\beta_{1},\beta_{2},\dots,\beta_{m})=\beta_{1}^{\prime},\beta_{2}^{\prime},\dots,\beta_{m}^{\prime}$. We summarize the algorithm in above theorem as follows.
Algorithm 2 finding an orthogonal basis
1:a lattice $\mathcal{L}=\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$
2:an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$
3:$\beta_{1}^{\prime},\beta_{2}^{\prime},\dots,\beta_{m}^{\prime}\Leftarrow\beta_{1},\beta_{2},\dots,\beta_{m}$
4:for $i=1$ to $m-1$ do
5:     $\beta_{i}^{\prime},\beta_{i+1}^{\prime},\dots,\beta_{m}^{\prime}\Leftarrow A(\beta_{i}^{\prime},\beta_{i+1}^{\prime},\dots,\beta_{m}^{\prime})$
6:end for
7:return $\beta_{1}^{\prime},\beta_{2}^{\prime},\dots,\beta_{m}^{\prime}$